2/5Farmer Reece took some bushels of corn to the market. She sold one-fourth of the bushels of corn to Farmer Peebles. She sold two-thirds of the bushels that were left to farmer Ricker. She sold one-sixth of the remaining bushels of corn to Farmer Condon. She sold three-fifths of the remaining bushels to Famer Skydancer Famer Reece brought the 20 bushels of corn that were left back to the farm. How many bushels of corn did Farmer Reece sell?
Answer by checkley71(8403)

(x*3/4)*1/3*5/6=20
(3x1*5)/(4*3*6)=20
15x/72=20
15x=20*72
15x=1440
x=1440/15
x=96 bushels had farmer Reece at the beginning.